A 1500kg wrecking ball hangs from a 20m -long cable. The ball is pulled back until the cable makes an angle of 30.0? with the vertical.

By how much has the gravitational potential energy of the ball changed?
Answer #1

delta U = m*g*h

= m*g*L*(1 - cos(30))

= 1500*9.8*20*(1 - cos(30))

= 39388.5 J
Answer #2

the height gain = 20 meters X sin30

PE = m g h
PE =1500 (9.8) x20sin30
EDIT

PE = 1500(9.8) x20 X 0.5
PE = 147000 Joules
Answer #3

Ok, so you will have to draw the cable making the angle of 30
Answer #4

chnge in gravitational potential energy = mgh

m = 1500kg

g = 9.8 m/s2

h = l -lcos 30 = 20(1-cos 30) =2.68m

change potential energy U = 1500x 9.8 x 2.68 =39396 J